According to a report* published in 1990 in the journal Anesthesia Progress, breathing nitrous oxide (N2O) causes “happy, euphoric mood changes ... and experiences of a dreamy, detached reverie state.”

The concentration used in the study was XN2O = 0.3. Our classroom is 3 m x 6 m x 6 m. If I bring a 5 lb (2.25 kg) tank of N2O to class, will it provide enough gas to induce euphoria? (you can estimate the temperature and pressure on a typical day)
Answer #1

Classroom can be considered as a cube and its volume V = a x b x c = 3 m x 6 m x 6 m = 108 m3

moles in 2.25 kg of N2O = 2.25 x 103 g / 44.013 g/mol = 51.1212596278 mol

1 m3 (cubic meter) = 103 L;

concentration = n/v = moles/liter = 51.1212596278 mol / (108 x 103 L) = 0.0005 M

XN2O concentration calculated = 0.0005 M, but observed euphoria concentration XN2O = 0.3

Answer: NO, it will not provide enough gas to induce euphoria.

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")